Chelsea Fashions is expected to pay an annual dividend of $1.20 a share next year. The market price of the stock is $27.00 and the growth rate is 4 percent. What is the firm's cost of equity?

8.49 percent

8.26 percent

8.44 percent

8.32 percent

8.13 percent
Answer #1

HI

Dividend next year D 1 = $1.20

Stock Price P = $27

growth rate g = 4%

cost of equity k =?

As per dividend discount model

Stock Price P = D1/(1k-g)

27 = 1.2/(k-0.04)

27k - 27*0.04 = 1.2

27k = 1.2 + 1.08

k 2.28/27 = 8.44%

Hence 8.44% is correct answer

Thanks